OpenStudy (anonymous):

lim x to 0=(x+xcosx)/(sinxcosx)

OpenStudy (anonymous):

this is what it lookslike in the book \[\frac{ \lim }{ x \rightarrow 0 }=\frac{ x+xcosx }{ sinxcosx }\]

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{x+x \cos(x)}{\sin(x)\cos(x)}=\lim_{x \rightarrow 0}\frac{x(1+\cos(x))}{\sin(x)\cos(x)}=\lim_{x \rightarrow 0}\frac{x}{\sin(x)} \cdot \lim_{x \rightarrow 0}\frac{1+\cos(x)}{\cos(x)}\] try this

OpenStudy (anonymous):

here is what i am thinking on this one, do the inverse of x/sin(x) so that it creates 1... times the 1+cosx/cosx

OpenStudy (anonymous):

and 1+cosx/x is 1+1/2

OpenStudy (freckles):

inverse?

OpenStudy (freckles):

x/sin(x) goes to 1 as x goes to 0

OpenStudy (freckles):

and all you have to do with the other function is direct substitution

OpenStudy (freckles):

so you would be correct to say (1+1)/1 is that what you meant?